OpenStudy (anonymous):

Someone please help me!! [9.02] What are the solutions of x^2+5x−24=0? Show all work for full credit.

OpenStudy (anonymous):

i dont understand how to do it

OpenStudy (butterflydreamer):

What two numbers can multiply together to give you -24 ? Then see if you can add or subtract these two numbers to give you +5

OpenStudy (anonymous):

-6x4

OpenStudy (anonymous):

Or you could use the quadratic formula to solve this quadratic equation

OpenStudy (anonymous):

solutions are: (x-3)(x+8)

OpenStudy (butterflydreamer):

ok. So can you add or subtract -6 and 4 to give you 5? If we try -6 + 4= -2 and if we try -6 - 4 = -10 So these two numbers don't work.

OpenStudy (anonymous):

radelp is right

OpenStudy (butterflydreamer):

Let's try -3 and 8. -3 x 8 = -24 so that works :) Now we try to add or subtract -3 and 8. So -3 + 8 = 5 So that means those are your answers and you would write it as (x-3) (x+8) = 0 Then solve for x

OpenStudy (butterflydreamer):

So now we want (x+3) = 0 This means x = -3 Because -3 + 3 =0 and then we do the same thing for (x+8) =0 So x=-8

OpenStudy (anonymous):

im still confused on how you got that...

OpenStudy (anonymous):

have you heard of the FOIL method?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

First outter inner last

OpenStudy (anonymous):

hu?

OpenStudy (butterflydreamer):